Ivanhoe Potions, Inc., a pharmaceutical company, bought a machine at a cost of $2 million five years ago that produces pain-reliever medicine. The machine has been depreciated over the past five years, and the current book value is $724,000. The company decides to sell the machine now at its market price of $1 million. The marginal tax rate is 35 percent.

What are the relevant cash flows?

How do they change if the market price of the machine is $600,000 instead?
Answer #1
after tax cashflow = selling price*(1-tax rate)+book value*tax rate =

selling price = 1000000

after tax cashflow = 1000000*(1-0.35)+724000*0.35=903400

selling price = 600000

after tax cashflow = 600000*(1-0.35)+724000*0.35=643400